Xentia Technologies Group (XTG) is considering investing in developing new 4D television technology. The CEO of XTG, Ms Jane Smith, has appointed you to evaluate the proposal for the board. If the new project goes ahead it is expected that it be operational at the beginning of year 2 (with the first revenues generated by the end of that year). Once the new project is operational it will render the company’s existing 2D technology project obsolete. The new project is then expected to have an operating life of six years.

To assist you in evaluating the project the following information has been prepared:

Existing 2D technology project:

Constant annual earnings before depreciation and taxes (EBDIT)            $400,000

Annual depreciation expense on equipment                                      $0

(Equipment fully depreciated)

Annual working capital balance                                                          $200,000

(Expected to last in perpetuity)

Expected salvage value of equipment if rendered obsolete              $0

New 4D technology project:

New equipment outlays (immediate)                                                  $10,000,000

Expected constant EBDIT                                                                  $3,800,000

(In the first year of operation)

Annual depreciation rate on equipment (straight line)                       10% p.a.

Expected salvage value of equipment at the end of the project            $3,500,000

Working capital requirement (once project is operational)                 $300,000

Additional Information:

  1. Company tax rate is 30%
  2. Xentia is financed with $25 million in market value of debt and $50 million in market value of equity.
  3. Xentia has a beta of 1.6.
  4. Revolutionary Technology Corporation (RTC) is currently using technology that is similar in risk profile to the new 4D project.
  5. RTC is financed with $40 million in market value of debt and $40 million in market value of equity.
  6. RTC has a beta of 1.75
  7. Xentia borrows debt capital at a cost of 8% p.a. compounded semi-annually
  8. The long term market risk premium (including franking credits) is 9.75% p.a.
  9. The current yield on Commonwealth Bonds is 4.25% p.a.
  10. Xentia operates in an imputation tax system.

Required:

  1. What is the appropriate discount rate that should be used to evaluate the project? Explain your decision.   

2. Calculate the NPV of the project. Present the cash flows used in the NPV calculation in a table.

3. Advise whether you should recommend the project. Explain your decision.
Answer #1
Ans 1.
From the RTC data we can find the ungeared beta =betaU
for the 4D industry that will give the indusry beta.
For RTC :
Value of Debt =D=$40M
Value of Equity =E=$40M
Tax rate T=30%
RTC beta =beta G=1.75
beta od Debt =betaD =assumed zero
So betaU= betaG*E/[E+D*(1-T)] +betaD*D*(1-T)/[E+D*(1-T)]
betaD=0
so betaU= 1.75*40/[40+0.7*40]
betU=1.03
So the ungeraed beta for 4D indutry =1.03
Using that to find the geared beta of Xentia Technologies
E =$50M =Equity value of Xentia
D =Debt value of Xential =$25M
Tax rate T=30%
betaU=1.03
Assume beta for the investment od Xentia =betaG
betaG=betaU+(betaU-betaD)*(1-T)*D/E
as betaD asssumed zero,
betaG=1.03+1.03*0.7*25/50
betaG=1.39
So we shall consider the Xentia beta for equity for the 4D
project as 1.39
Risk free rate =4.25%
Market Risk premium ( with franking credit)=9.75%
So Cost of Equity =Re=Rf+betaG*Rp
Re=4.25%+1.39*9.75%
Re=17.8%
So cost of equity for Xenita =17.8%
Cost of Capital for Xenita debt =8% pa compounded semi annually
EAR =(1+8%/2)^2-1=8.16% pa
Post Tax cost of Debt =8.16%*(1-30%)=5.71%
Calculating WACC
Capital Type Market Value $M Weightage on Market Value Post TAx cost Weighted cost
Debt 25 33% 5.71% 1.90%
Equity 50 67% 17.80% 11.87%
Total 75 100% 13.77%
So WACC = Discount rate for the project =13.77%
New Eqp cost =$10M
SL depreciation is 10% , so machine life in 10 years, salvage at the end of 10 yrs not given ,
so assuming annual depreciation $$1,000,000
Book value after 6 yrs            4,000,000
Salvage value after 6 yrs            3,500,000
Capital Loss               500,000
Tax saves on Capital Loss @30%=                       150
Total effective salvage cash flow with tax savings            3,500,150
NPV Caluclation Amts in $
Particulars Year 0 Year 1 Year 2 Year 3 Year 4 Year 5 Year 6
Initial Investment
New Equipment        (10,000,000)
Incremental Working capital Required ( assuming to be arranged at the end of year 1 , as operations starts in the beginning of yr 2)           (100,000)
Total Initial Investment        (10,000,000)           (100,000)
Cash flow from Operations
Incremental EBDIT ($3,800,000-$400,000)                3,400,000          3,400,000           3,400,000                  3,400,000           3,400,000
Less Incremental Depreciation         1,000,000                1,000,000          1,000,000           1,000,000                  1,000,000           1,000,000
EBT        (1,000,000)                2,400,000          2,400,000           2,400,000                  2,400,000           2,400,000
Tax @30%           (300,000)                   720,000             720,000              720,000                     720,000               720,000
PAT           (700,000)                1,680,000          1,680,000           1,680,000                  1,680,000           1,680,000
Add Back depreciation         1,000,000                1,000,000          1,000,000           1,000,000                  1,000,000           1,000,000
Net Operational cash flow             300,000                2,680,000          2,680,000           2,680,000                  2,680,000           2,680,000
Terminal Cash flows
After Tax Salvage value (with Tax saved )           3,500,150
Return of Incremental WC               100,000
Total Terminal Cash flow           3,600,150
Total Cash flow=a+b+c=        (10,000,000)             200,000                2,680,000          2,680,000           2,680,000                  2,680,000           6,280,150
PV factor @13.77%=1/1.1377^n 1